OpenStudy (anonymous):

solve for x 5(2x-1)=6

OpenStudy (anonymous):

Whats the answer to my question Please tell me

OpenStudy (anonymous):

1x i believe

OpenStudy (anonymous):

5(2x-1)-6=0 10x-11=0 10x=11 x=11/10